### v1

# Controlled Low-Rank Adaptation with Subspace Regularization for Continued Training on Large Language Models

Abstract
Large language models (LLMs) exhibit remarkable capabilities in natural language processing but face catastrophic forgetting when learning new tasks, where adaptation to a new domain leads to a substantial decline in performance on previous tasks. In this paper, we propose Controlled LoRA (CLoRA), a subspace regularization method on LoRA structure. Aiming to reduce the scale of output change while introduce minimal constraint on model capacity, CLoRA imposes constraint on the direction of updating matrix’s null space. Experimental results on commonly used LLM finetuning tasks reveal that CLoRA significantly outperforms existing LoRA subsequent methods on both in-domain and out-domain evaluations, highlighting the superority of CLoRA as a effective parameter-efficient finetuning method with catastrophic forgetting mitigating. Further investigation for model parameters indicates that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
Controlled Low-Rank Adaptation with Subspace Regularization for Continued Training on Large Language Models
Yuheng Lu1, Bingshuo Qian1, Caixia Yuan1, Huixing Jiang2, Xiaojie Wang1,thanks: Corresponding author 1Beijing University of Posts and Telecommunications, 2LI Auto Inc. {yuheng.lu, bsqian, yuancx, xjwang}@bupt.edu.cn, jianghuixing@lixiang.com
1 Introduction
Large language models (LLMs) demonstrate remarkable capabilities in natural language tasks. However, when performing continued training on additional datasets, a key challenge may faced, known as catastrophic forgetting McCloskey and Cohen (1989), where adaptation to a new domain leads to a substantial decline in performance on previous tasks.
Existing approaches to mitigate catastrophic forgetting can be broadly categorized into data-based, architecture-based, and learning-based methods Wang et al. (2023a). Data-based methods de Masson D’Autume et al. (2019) primarily based on rehearsing prior training data, which raises data privacy concerns. Additionally, for LLMs, obtaining the necessary prior training data samples is challenging due to their training on massive datas. Architecture-based methods Wang et al. (2023c); Razdaibiedina et al. (2023) introduce isolated parameters for each continued training stage for reducing interference. In contrast, learning-based methods train in the shared vector space, controlling learning process by adding regularization terms to the loss or employing specific optimization designs. Inference for architecure-based methods typically involves a selection process Gurbuz and Dovrolis (2022); Kang et al. (2022), which is more complex than that for learning-based methods. As continued trained LLMs are generally regarded as foundation models, flexibility is essential for their broader applications. Consequently, due to deployment considerations, learning-based methods are preferred over architecture-based methods for LLMs.
The core idea of learning-based methods is to constrain parameter updates, which aligns precisely with the Parameter-Efficient FineTuning (PEFT) research paradigm of LLMs. Although initially proposed for computational efficiency, PEFTs have demonstrated to learn less and forget lessBiderman et al. (2024), primarily due to their constrained model capacity. Notably, a well-established insight that related to learning-based methods in PEFT research is that LLMs are primarily finetuned within a specific low-rank subspace, this insight has led to the development of the Low-Rank Adaptation method (LoRA)Hu et al. (2021).
However, LoRA imposes no restrictions on parameter updates beyond the low-rank constraint, and matrix perturbation theory suggests that even low-rank updates can significantly influence matrix properties Sherman (1949); Davis and Kahan (1970). For instance, in an extreme case, it is theoretically possible to learn a model that eliminates all top-k principal components (optimal rank-k approximation) through a rank-k update, thus destroy most of the base model’s ability. Therefore, LoRA would be benifit from more constraints for mitigating catastrophic forgetting. However, more constraints would reduce model capacity for updating, which influences the effectiveness of training. For instance, adding L2 regularization significantly restricts the norm of the updating matrix. Consequently, effective management of the capacity-forgetting balancing has become a major concern.
[…]
Figure 1: Illustration of the intuition behind our approach. For input $x$, the component in $\mathrm{Null}(\Delta W)$ (null space of the updating matrix $\Delta W$) would be ignored, the change of output $\Delta y$ is obtained only from the component in $\mathrm{Row}(\Delta W)$ (row space of $\Delta W$, the orthogonal complement of $\mathrm{Null}(\Delta W)$). CLoRA introduces a pre-defined subset of $\mathrm{Null}(\Delta W)$ by imposing orthogonal regularization with pre-defined matrix $P$.
To address this concern, in this work, we propose Controlled LoRA (CLoRA), a subspace regularization method on LoRA structure. We start the design of CLoRA from the perspective of the null space of updating matrix. The intuition behind CLoRA is illustrated in Figure 1, where the output change $\Delta y$ is derived from applying the updating matrix $\Delta W$ on the component of the input $x$ that falls within the row space of $\Delta W$, while components in the null space are ignored. Under this intuition, for reducing the scale of output change, options include reducing the scale of $\Delta W$, and encouraging more input component fall in the null space of $\Delta W$. The former is more related to model capacity, and for concerns of capacity-forgetting balancing, we focus on the latter.
The dimension of the null space for the updating matrix is directly determined by the rank of it, LoRA already addressed. A key factor remains, the direction of null space, which influence input components that fall in, but free-learned LoRA does not constraint. CLoRA constraint the direction of null space of updating matrix by introducing a pre-defined subset of that, this is implemented by orthogonal regularization with a pre-defined matrix. Unlike methods that impose restrictions on rank or norm, which significantly influence model capacity, CLoRA introduces constraint on the direction of the null space. Experimental results on commonly used LLM finetuning evaluations demonstrate that our proposed CLoRA outperforms existing methods on both in-domain downstream tasks and out-domain evaluations. This indicates the superiority of CLoRA as an effective approach for parameter-efficient finetuning with catastrophic forgetting mitigating. Additionally, we take analysis on parameters of the learned model, results show that CLoRA reduces the scale of output change with minimal impact on model capacity.
Our contributions are summarized as follows,
  • We propose CLoRA, a subspace regularization method on LoRA structure, which serves as an advanced parameter-efficient finetuning technique with catastrophic forgetting mitigating for LLMs.
  • Our proposed CLoRA demonstrates superior performance on commonly used LLM finetuning tasks compared to existing methods. It has also shown remarkable mitigating of catastrophic forgetting in out-domain evaluations.
  • Parameter investigation results indicate that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
2 Related Works
2.1 Reducing Catastrophic Forgeting
Catastrophic forgetting is a significant challenge in various transfer learning scenarios, including continual learning Wang et al. (2023a) and LLM continued training Wu et al. (2024). In these settings, continued training on new tasks may impair abilities of the pre-trained model. Approaches for mitigating catastrophic forgetting can be broadly categorized into data-based, architecture-based and learning-based methods.
Data-based methods
primarily based on rehearsal of prior training data or representation, de Masson D’Autume et al. (2019) introduce an episodic memory for experience rehearsal, Rebuffi et al. (2017); Chaudhry et al. (2019) selects previous training data for rehearsaling. For LLMs, acquiring the necessary prior training data is challenging due to the extensive amount of data used in their training. Instead, the concept of rehearsal is commonly adopted by mixing data from general domains for LLM continued training. This approach is generally orthogonal to model-related methods, thus we will not discuss it further.
Architecture-based methods
Wang et al. (2023c); Razdaibiedina et al. (2023) introduce isolated parameters for each continued training stage to reduce interference. Wang et al. (2023c) use isolated parameters for each task, and enables a selecting mechenism during inference. Progressive Prompts Razdaibiedina et al. (2023) sequentially concatenates prompts for each task with previously learned prompts. These architecture-based methods generally require specific techniques for inference and continued training, resulting in a lack of flexibility, particularly in the context of LLMs.
Learning-based methods
performs continued training in a shared vector space, controlling learning process by adding regularization term on loss or applying specific optimization designs. Orthogonal Gradient Descent (OGD) Farajtabar et al. (2020) constrains the parameters to move within the orthogonal space defined by the gradients of previous tasks. Continual LoRA (C-LoRA)Smith et al. (2024) and O-LoRA Wang et al. (2023b) introduce regularization with previous continual learned parameters for reducing interference in the multi-stage training setting. Our proposed CLoRA imposes orthogonal regularization similar to O-LoRA, but the regularization matrix is not restricted to be the previous learned parameter, thus CLoRA can be used for one-stage continued training whereas O-LoRA cannot.
2.2 LoRA and Subspace Tuning
Parameter-Efficient FineTuning (PEFT) Han et al. (2024) aims to tune models with minimizing computational resources, which is widely used for large-scale models including LLMs. Among these methods, LoRA Hu et al. (2021) and its subsequent variants Wang et al. (2024a); Liu et al. (2024) learn a low-rank decomposition for updating parameter matrices, and could be categorized into learning-based continued training method, which is the focus of our work.
The core insight of LoRA is to tune model within a low-rank subspace, and with no additional constraints imposed on this tuning subspace. Some subsequent works delve deeper into the tuning subspace to mitigate catastrophic forgetting for LLM continued training, MiLoRA Wang et al. (2024a) and PiSSA Meng et al. (2024) use singular value decomposition (SVD) components of the original parameters for LoRA initialization, with MiLoRA uses minor components while PiSSA uses major components; O-LoRA Wang et al. (2023b) introduce orthogonal regularization for each LoRA subspace. Our proposed CLoRA also falls within this category, differing from the selection and utilization of the focused subspace.
[…]
Figure 2: Illustration of CLoRA on typical decoder-only transformer based LLMs. LoRA updating is applied on v-proj in multi-head attention layer for each layer. CLoRA add orthogonal loss computes from trainable LoRA parameters (A and B) to the original language modeling loss.
3 Prelimilaries
Notation Description
$W$ parameter matrix in base model
$\Delta W$ updating of the parameter
$x$ input for $W$
$y$ output for $W$, $y=Wx$
$\Delta y$ output change, $\Delta y=\Delta Wx$
$||v||$ L2 norm of vector $v$
$||A||$ L2 norm(largest singular value)
of matrix $A$
$r$ rank of updating matrix
$k$ number of regularization vectors
Table 1: Notations.
3.1 Notations
The notations commonly used in this paper are summarized in table 1. We provide some additional notes here. While generally used for denote the input and output of the while model, we denote $x,y$ as input and output to a single linear layer, represented by $W$. $||A||$ denotes L2 norm (largest singular value) in our paper, instead of Frobenius norm ($||A||_{F}=\sqrt{\sum A[i,j]^{2}}$). $r$ and $k$ are most important hyperparameters for CLoRA, $r$ is the rank of updating matrix, which is used in all LoRA works, $k$ is the number of regularization vectors(column of regularization matrix) in CLoRA.
3.2 Problem Definition
We aim to mitigate catastrophic forgetting in LLM continued training. Catastrophic forgetting menifest as performance decline on tasks from previous domain when training on new domain. To evaluate this, we conduct both in-domain tasks (demonstrating the effectiveness of training) and out-domain tasks (from previous domain, indicating the degree of forgetting). We conduct experiments in the one-stage tuning setting, specifically, we finetune a base LLM on one training dataset, then take in-domain and out-domain evaluations. Note that there is no clear domain specific for base LLMs, but benchmarks exist for evaluating the ability of LLMs on wide range of domains Gao et al. (2024), and we take those with minimal overlap with training data for out-domain evaluation.
4 Method
In this section, we introduce Controlled Low-Rank Adaptation (CLoRA) method. We illustrate the application of CLoRA in transformer-based LLMs in Figure 2. CLoRA shares the same modeling structure with LoRA, but imposes on orthogonal regularization term computed using LoRA parameters into the loss function.
CLoRA Modeling
Consistent with LoRA, CLoRA decomposes the updating for a parameter matrix $W$ to a multiplication of two lor-rank matrices $\Delta W=AB^{T}$, where $W,\Delta W\in\mathbb{R}^{m\times n}$, $A\in\mathbb{R}^{m\times r}$, $B\in\mathbb{R}^{n\times r}$, $r\ll m,n$.
CLoRA computes orthogonal regularization for $A$ and $B^{T}$ with untrainable pre-defined matrix $P_{A}\in\mathbb{R}^{m\times k}$ and $P_{B}\in\mathbb{R}^{n\times k}$, where $k$ is a hyperparameter controlling the size of regularization matrix, larger $k$ introduces more constraint. The orthogonal regularization loss on one LoRA parameter $A$ is defined as
$L_{orth}(A,P_{A})=\sum_{i,j}||AP_{A}^{T}[i,j]||^{2}$ (1)
where $A\in\mathbb{R}^{m\times r}$, $P_{A}\in\mathbb{R}^{m\times k}$. $L_{orth}(A,P_{A})$ regularize on orthogonality of every $(A[:,i],P_{A}[:,j])$ pairs. The final loss of CLoRA in a transformer-based LLM is defined as
$L_{LM}(\Theta,input)+$ (2)
$\lambda\sum_{i}(L_{orth}(A_{i},P_{A_{i}})+L_{orth}(B_{i}^{T},P_{B_{i}}^{T}))$
where $L_{LM}(\Theta,input)$ is the original language model loss on text input and LLM parameters $\Theta$, the summation on $L_{orth}$ is over index of all trainable parameter matrices. $\lambda$ controls the weighting of orthogonal loss, we set it to 1 as default.
Initialization
Following LoRAHu et al. (2021), we initialize $A$ with gaussian noise and $B$ is zeros, ensuring $\Delta W$ is zero at the begining of training.
For the CLoRA regularization matrices, we suggest that they should have othogomal rows. Specifically, for regularization matrix $P\in\mathbb{R}^{m\times k}$, $||P[:,i]||=1$ for every $i$, and $P[:,i]P[:,j]=0$ for $i\neq j$. This ensures uniform regularization. We propose several variants as follows,
  • Random Initialize: Initialize with an orthogonal matrixSaxe et al. (2014).
  • Initialize from SVD: Previous works Sharma et al. (2024); Wang et al. (2024a) have explored the roles of singular value decomposition (SVD) components in LLM parameters. We adopt this intuition to initialize CLoRA regularization matrices from SVD. For a SVD decomposition of parameter $W=USV^{T}$ with rank $r$, where $W\in\mathbb{R}^{m\times n}$, $U\in\mathbb{R}^{m\times r}$, $S\in\mathbb{R}^{r\times r}$ is a diagonal matrix, $V\in\mathbb{R}^{n\times r}$. For CLoRA updating $\Delta W=AB^{T}$, we initialize the regularization matrix $P_{A}\in\mathbb{R}^{m\times k}$ as $U[:,\mathbf{s}]$, $P_{B}\in\mathbb{R}^{n\times k}$ as $V[:,\mathbf{s}]$, where $\mathbf{s}$ is a list of selecting index with length $k$.
5 Experiment and Analysis
In this section, we conduct experiments to evaluate and analyze our proposed CLoRA method. All evaluations are available in lm-eval Gao et al. (2024) with MIT License. We aim to answer these research questions,
  • RQ1: Does CLoRA perform effectively as a parameter-efficient finetuning method with catastrophic forgetting mitigating?
  • RQ2: How do the size and selection of regularization matrices influence the performance of CLoRA?
  • RQ3: How does CLoRA demonstrate superiority on capability-forgetting balancing?
5.1 Evaluating for Tuning on Commonsense Reasoning Dataset (RQ1)
To answer RQ1, we conduct experiments following the setting in the previous works on PEFTLiu et al. (2024); Wang et al. (2024a). Specifically, we finetune LLaMA-2-7B Touvron et al. (2023) with a dataset for commonsense reasoning. We evaluated the model on both in-domain downstream tasks to evaluate learning effectiveness and on out-domain tasks with minimal overlap with the training dataset to measure the degree of forgetting. We report partial results here, full results on each CLoRA variants and results on finetuning LLaMA-3-8BAI@Meta (2024) are reported in Appendix B.
5.1.1 Experiment Setup
Datasets and Tasks
We use Commonsense170K Hu et al. (2023) for finetuning. For in-domain evaluation, eight commonsense reasoning datasets are used, including BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag(HS) Zellers et al. (2019), WinoGrande(WG) Sakaguchi et al. (2019), ARC-e, ARC-c Clark et al. (2018), and OBQA Mihaylov et al. (2018). The tasks are formulated as multiple-choice problem, and we report accuracy based on the last checkpoint. For out-domain evaluations, BIG-Bench-Hard(BBH) Suzgun et al. (2022) and MMLU-Pro Wang et al. (2024b) are used. These benchmarks encompass challenging subsets of tasks across a wide range of domains and are widely employed for evaluating the capabilities of LLMs. Additionally, they include samples that are more complex than those in our training data, ensuring minimal overlap. We use lm-eval Gao et al. (2024) for reporting out-domain evaluation.
Comparison Methods
  • LoRA Hu et al. (2021) is a widely-used parameter-efficient finetuning technique, and it serves as the foundation of our proposed CLoRA.
  • DoRA Liu et al. (2024) is a recent work on structure improvement of LoRA, we include it as a baseline for improved LoRA.
  • PiSSA Meng et al. (2024) and MiLoRA Wang et al. (2024a) are two variants of LoRA, both employing SVD components for LoRA initialization, MiLoRA use minor components while PiSSA use major. Notably, MiLoRA can be categorized as a catastrophic forgetting mitigating method.
  • Reducing the updating rank(-r*): Lower rank $r$ imposes stricter constraints on the updating matrix. We maintain a consistent rank across all methods and consider variations in rank as a separate baseline.
  • L2 regularization(-L2) introduces L2 regularization for trainable parameters, serving as a fundamental approach to limit updates.
  • CLoRA: Random initialize variant of CLoRA, it is the basic setting of our method. We will include results for other variants in the subsequent section 5.2.
Category Method BoolQ PIQA SIQA HS WG ARC-e ARC-c OBQA Avg.
ChatGPT ChatGPT 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LoRA 71.9 80.9 78.9 90.3 83.5 83.0 70.2 80.8 79.9
LoRAs DoRA 73.0 81.9 80.3 90.2 82.8 84.6 69.4 81.8 80.5
PiSSA 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA 71.5 82.0 80.0 91.0 83.0 82.3 68.9 81.2 80.0
Reducing LoRA-r8 71.0 80.5 78.1 90.0 83.0 81.1 68.5 78.0 78.8
Forgetting LoRA-r16 71.0 81.8 78.9 90.3 81.1 83.1 69.7 82.2 79.8
LoRA-L2 70.3 83.0 80.2 92.7 83.1 84.2 71.2 81.4 80.8
Ours CLoRA 73.7 84.5 80.9 94.5 85.9 88.1 75.9 86.0 83.7
Table 2: In-domain results on commonsense reasoning evaluations, with accuracy scores (%) reported. Bold font indicates the highest performance for each dataset across the different PEFT methods.
Hyperparameter
We use the same hyperparameter configurations as Wang et al. (2024a). Details are listed in Appendix A. Notably, we use 32 for updating matrix rank $r$ as default for all methods. For CLoRA regularization matrix, we select $k$ in $[128,256,512,1024,2048]$. For LoRA-L2, 1e-5 is used for weighting of L2 regularization, we note that 1e-4 is also tested, but too large for getting effective tuning.
5.1.2 Results and Analysis
In-domain Evaluation
We report in-domain results in Table 2. LLaMA-2-7B cannot be directly applied to these tasks without finetuning, following previous worksWang et al. (2024a), we include the ChatGPT baseline reported in Liu et al. (2024) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought Wei et al. (2022). For CLoRA, we report the best result across all results for all $k$.
For in-domain evaluation, results demonstrate that CLoRA outperforms on all datasets, surpassing the best baseline by 2.9 points in average accuracy (%). These outcomes suggest that, although proposed primarily for mitigating catastrophic forgetting, CLoRA also performs as a effective PEFT method. We attribute this to the nature of LLM finetuning, as an instance of transfer learning, the performance of LLM finetuning is strongly correlated with the base model’s ability, when catastrophic forgetting occurs during training, the base model’s strength may diminish. Therefore, we hypothesize that a method with effective capacity-forgetting balancing would exhibit strong effectiveness in LLM finetuning.
Method BBH MMLU-Pro
base 34.91 18.56
PiSSA 29.54 11.33
DoRA 28.24 11.67
LoRA 26.69 14.46
MiLoRA 25.14 17.74
LoRA-r8 26.90 14.58
LoRA-r16 26.73 11.54
LoRA-L2 32.93 16.59
CLoRA 38.67 20.59
Table 3: Out-domain results on LLM benchmarks, with accuracy scores (%) reported.
12825651210242048767676767878787880808080828282828484848486868686kAvg. Accuracy (%)CLoRA-randomCLoRA-majorCLoRA-minor
1282565121024204820202020222222222424242426262626282828283030303032323232kAvg. Accuracy (%)CLoRA-randomCLoRA-majorCLoRA-minor
Figure 3: Performace for CLoRA on commonsence reasoning finetuning setting in-domain(above), out-domain(below) with different variant and k.
Out-domain Evaluation
We report out-domain results for CLoRA and baselines in Table 3. “base” denotes the result on the base model LLaMA-2-7B, “CLoRA” denotes the result of best model from in-domain evaluation.
In out-domain evaluation, all baselines underperform compared to the base model, highlighting the severe issue of catastrophic forgetting in this experimental setup. Notably, our proposed CLoRA not only outperforms all baselines, but also surpasses the base model’s performance. We attribute this to the effective capacity-forgetting balancing for CLoRA, enabling the effective extraction of generally useful knowledge from the commonsense reasoning training dataset.
Answer for RQ1:
The superior performance in both in-domain and out-domain evaluations demonstrates that our proposed CLoRA serves effectively as a parameter-efficient finetuning method with catastrophic forgetting mitigating.
5.2 Evaluating for different CLoRA variants and hyperparameters (RQ2)
To answer RQ2, we compare results for various CLoRA variants to demonstrate the impact of regularization matrix’s size and selection.
Datasets and Tasks
Same as in section 5.1, we finetune LLaMA-2-7B on Commonsense170K, and conduct in-domain evaluation on eight commonsense reasoning tasks and out-domain evaluation on two LLM benchmarks. We report average results for each evaluation. To ascertain whether the optimal $k$ value is consistent across tasks, we additionally perform an experiment in the math domain. we finetune LLaMA-2-7B on the MetaMathQA Yu et al. (2024), which contains 395K samples augmented from the training set of GSM8K Cobbe et al. (2021) and MATHHendrycks et al. (2021). We use test set of GSM8K and MATH for evaluation and report the results on the last checkpoint.
Comparision Methods
  • Different $k$(-k*): We test CLoRA with different regularization matrix size ($k$). Higher $k$ imposes stricter constraints on the updating matrix.
  • CLoRA-random: CLoRA with random initialization.
  • CLoRA-major: Use SVD major components to initialize CLoRA regularization matrix.
  • CLoRA-minor: Use SVD minor components to initialize CLoRA regularization matrix.
For math evaluation, PiSSAMeng et al. (2024) and MiLoRAWang et al. (2024a) are included as baselines, we use CLoRA-major for CLoRA variant.
5.2.1 Results and Analysis
Commonsense Reasoning Setting
We illustrate the performance for in-domain and out-domain evaluations with different $k$ for commonsense reasoning setting as in Figure 3. The results indicate that for all CLoRA variants tested, as $k$ increases, the scores of both in-domain and out-domain evaluation generally exhibit an upward trend. Despite similar performance across variants under the same $k$, CLoRA-major generally performs better, highlighting the advantage of preserving the well-learned major components Wang et al. (2024a) of the base model.
Math Setting
Results on math setting are reported in Table 4. CLoRA outperforms other baselines, and the optimal $k$ value found to be around 128, differing from the 2048 in commonsense reasoning setting. This discrepancy can be attributed to the complexity of math tasks, which necessitate greater model capacity during finetuning. These findings suggest that adjusting $k$ benefits the balancing of capacity and forgetting, demonstrating the flexibility of our approach.
Method GSM8K MATH
LoRA 60.58 16.88
PiSSA 58.23 15.84
MiLoRA 63.53 17.76
CLoRA-k64 64.29 17.52
CLoRA-k128 64.59 18.38
CLoRA-k256 63.45 17.58
Table 4: Math evaluation on GSM8K and MATH, with accuracy scores (%) reported.
Answer for RQ2:
Results indicate that the choice of regularization matrix does influence the effectiveness of CLoRA, albeit not significantly. Generally, we recommend using random initialization (CLoRA-random) or initialization from major SVD components (CLoRA-major). Regarding the choice of $k$, as k increases, a trend of performance improvement followed by a decline is observed. Different tasks may correspond to different optimal $k$, we suggest that a smaller $k$ could be used for more challenging tasks.
5.3 Understanding Capacity-Forgetting Balancing(RQ3)
To answer RQ3, we investigate the parameter of trained models to quantify the capacity-forgetting balancing issue.
Measuring Model Capacity and Degree of Forgetting
Consider that catastrophic forgetting primarily arises from output changes caused by parameter updating, the greater the impact of these updates, the more severe the catastrophic forgetting may be. We measure the degree of forgetting with the relative scale of output change in the parameter level, to be specific, for updating matrix $\Delta W$, with input $x$, the relative scale of output change (denoted as $\mathrm{F}$) is defined as
$\mathrm{F}(\Delta W,x)=\frac{||\Delta Wx||}{||x||}$ (3)
We highlight the role of $x$ in the measurement of “F”, as it reflects the real world use case. Specifically, we sample 100 text data from test set and collect input $x$ for each parameter from the model forward pass.
To measure model capacity, we note that there is a gap between theoretical capacity of a model Abu-Mostafa (1989) and the practical outcome of the learned model. Therefore, we delegate the measurement of model capacity to the scale of the parameters in the learned model. Specifically, we measure the L2 norm $||\Delta W||$ for each updating parameter matrix.
Results and Analysis
We report the measurements averaged over all tokens and all updating parameters in Table 5. All models use LoRA rank $r$ of 32 unless specified otherwise.
The “reference” row is computed using the LoRA trained model, noting the output scale of original parameter $W$ instead of $\Delta W$. Compared with “reference” and “LoRA”, the difference of “F” is not very far, suggesting that LoRA training indeed introduces significant output change, thus still prone to catastrophic forgetting.
For “MiLoRA”, although intuitively promising, without effective control during training, it did not mitigate catastrophic forgetting, as evidenced by both downstream evaluations (in section 5.1) and the similar “F” and $\Delta W$ with LoRA.
For LoRA with lower rank (r8 and r16), after training, with $||\Delta W||$ indicates the reduction of capacity, the scale of output change does not show a decrease. Although theoretically, reducing the rank of the update matrix can increase the dimension of the null space and help to reduce the scale of output change, results not show this case. This suggests that altering $r$ may not a effctive way to alter forgetting.
For LoRA-L2, “F” indicates that it indeed mitigate forgetting, but in a large cost of capacity, demonstrated by the very small $||\Delta W||$.
For our proposed CLoRA, “F” shows a significantly reduce the scale of output change, while a relatively larger $||\Delta W||$ is maintained. This indicates that CLoRA minimizes catastrophic forgetting caused by large updates while having a subtle impact on model capacity. Thus we answer RQ3 that CLoRA performs effectively on capacity-forgetting balancing.
Method $||\Delta W||$ $\mathrm{F}$
reference 2.42
LoRA 22.63 0.79
MiLoRA 24.32 0.92
LoRA-r16 12.70 1.03
LoRA-r8 6.45 0.95
LoRA-L2 2.07 0.29
CLoRA-k128 10.84 0.36
CLoRA-k256 10.25 0.34
CLoRA-k512 8.19 0.27
CLoRA-k1024 6.64 0.21
CLoRA-k2048 5.00 0.14
Table 5: Measuing of model updating capacity($||\Delta W||$, larger for more capacity) and degree of forgetting($\mathrm{F}$, lower for less forgetting) for trained models.
6 Conclusion
In this paper, we introduce Controlled Low-Rank Adaptation(CLoRA), a simple yet effective parameter-efficient finetuning method for LLMs that mitigates catastrophic forgetting. We investigate the effectiveness of CLoRA on commonsense reasoning and math reasoning tasks. Experiment results demonstrate that CLoRA consistently outperforms previous LoRA-based methods on in-domain evaluations, highlighting the effectiveness of CLoRA as a parameter-efficient finetuning method. Additionally, out-domain evaluations demonstrate that CLoRA exhibits strong capabilities in mitigating catastrophic forgetting. Further investigation for model parameters indicates that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
7 Limitations
There are still several limitations that we reserve for future work: 1) While we have noted that our proposed CLoRA would benefit from the dedicated selection of regularization matrix to retain specific model abilities, we did not go further into this aspect. In this paper, we only explored the SVD initialization variant alongside the basic random initialization. 2) We delegate the measurement of model capacity and degree of forgetting to simple measurement of scale. Although these measurements reveal significant differences between CLoRA and previous works, we believe that further investigation would aid in the design of methods with stronger capacity-forgetting balancing capability.
Appendix A Detailed Experiment Setups
A.1 Hyperparameter Settings
Table 6 shows our detailed hyperparameters. This setting follows MiLoRAWang et al. (2024a) and DoRALiu et al. (2024).
A.2 Computation Environment
All of our experiments are conducted on 8 NVIDIA A800 GPUs. All methods for LoRA subsequents use Huggingface peft library [^1] https://github.com/huggingface/peft, training is conducted using trainer in Huggingface transformers library [^2] https://github.com/huggingface/transformers, with DeepSpeed ZeRORajbhandari et al. (2020) intergration.
Appendix B Full Experiment Results
B.1 Full Results on Commonsense Finetuning
Model PEFT BoolQ PIQA SIQA HS WG ARC-e ARC-c OBQA Avg.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA-2-7B LoRA 71.9 80.9 78.9 90.3 83.5 83.0 70.2 80.8 79.9
PiSSA 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA 71.5 82.0 80.0 91.0 83.0 82.3 68.9 81.2 80.0
DoRA 73.0 81.9 80.3 90.2 82.8 84.6 69.4 81.8 80.5
LoRA-r8 71.0 80.5 78.1 90.0 83.0 81.1 68.5 78.0 78.8
LoRA-r16 71.0 81.8 78.9 90.3 81.1 83.1 69.7 82.2 79.8
LoRA-L2-0.0001 - - - - - - - - -
LoRA-L2-0.00001 70.3 83.0 80.2 92.7 83.1 84.2 71.2 81.4 80.8
CLoRA-random-k128 72.7 84.1 77.7 91.6 83.0 85.3 69.9 81.6 80.7
CLoRA-random-k256 71.3 83.2 79.1 92.4 83.2 84.5 71.0 81.0 80.7
CLoRA-random-k512 72.8 83.0 79.5 93.0 83.9 85.7 73.0 84.8 82.0
CLoRA-random-k1024 73.3 84.8 79.6 91.1 86.1 86.9 73.1 85.6 82.6
CLoRA-random-k2048 73.7 84.5 80.9 94.5 85.9 88.1 75.9 86.0 83.7
CLoRA-major-k128 72.4 81.9 77.9 83.9 82.4 84.4 70.0 82.6 79.4
CLoRA-major-k256 73.2 83.5 79.6 93.0 83.3 88.1 72.6 84.2 82.2
CLoRA-major-k512 73.6 83.7 79.9 93.4 83.9 86.4 73.0 86.0 82.5
CLoRA-major-k1024 73.2 85.5 80.5 94.3 85.7 87.2 75.9 85.4 83.5
CLoRA-major-k2048 73.9 84.8 80.6 95.0 85.3 87.7 76.5 84.6 83.6
CLoRA-minor-k128 71.5 82.7 78.7 91.8 83.2 85.0 70.9 81.6 80.7
CLoRA-minor-k256 72.6 83.5 80.2 91.3 85.4 85.4 72.1 83.6 81.8
CLoRA-minor-k512 73.0 84.0 80.1 93.1 82.0 86.4 72.9 84.4 82.0
CLoRA-minor-k1024 73.1 83.7 79.2 93.7 84.8 87.1 73.2 83.2 82.3
CLoRA-minor-k2048 72.9 84.2 80.8 93.7 85.3 87.2 73.5 86.0 83.0
LLaMA-3-8B LoRA 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
PiSSA 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
DoRA 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8 85.2
CLoRA-random-k128 75.5 89.1 81.6 95.9 87.9 92.6 81.8 86.8 86.4
CLoRA-random-k256 75.3 88.8 81.4 85.7 88.7 92.7 82.3 88.4 85.4
CLoRA-random-k512 75.9 89.3 82.6 96.3 88.9 92.1 82.9 86.8 86.9
CLoRA-random-k1024 76.5 89.1 82.1 96.3 88.6 93.0 81.7 90.0 87.2
CLoRA-random-k2048 76.2 90.0 82.7 96.6 88.8 93.3 83.4 89.2 87.5
Table 7: In-domain results on commonsense reasoning evaluations, with accuracy scores (%) reported. Bold font indicates the highest performance for each dataset across the different PEFT methods for each base model.
Model PEFT BBH MMLU-Pro Avg.
LLaMA-2-7B - 34.91 18.56 26.74
LoRA 26.69 14.46 20.58
PiSSA 29.54 11.33 20.44
MiLoRA 25.14 17.74 21.44
DoRA 28.24 11.67 19.96
LoRA-r8 26.90 14.58 20.74
LoRA-r16 26.73 11.54 19.13
LoRA-L2-0.00001 32.93 16.59 24.76
CLoRA-random-k128 30.82 12.07 21.45
CLoRA-random-k256 31.92 17.81 24.87
CLoRA-random-k512 34.32 17.00 25.66
CLoRA-random-k1024 36.49 19.52 28.01
CLoRA-random-k2048 38.67 20.59 29.63
CLoRA-major-k128 32.69 18.09 25.39
CLoRA-major-k256 35.11 18.89 27.00
CLoRA-major-k512 35.81 19.88 27.85
CLoRA-major-k1024 37.06 19.73 28.40
CLoRA-major-k2048 38.83 20.08 29.46
CLoRA-minor-k128 34.06 17.03 25.55
CLoRA-minor-k256 33.16 17.11 25.13
CLoRA-minor-k512 35.42 18.97 27.20
CLoRA-minor-k1024 37.08 18.87 27.98
CLoRA-minor-k2048 40.96 20.37 30.67
Table 8: Out-domain results on two LLM benchmarks, with accuracy scores (%) reported. Bold font indicates the highest performance for each benchmark across all methods.
We report the full results that we conducted on in-domain evaluation(Table 7) and out-domain evaluation(Table 8) for commonsense reasoning finetuning. All models use LoRA rank $r$ of 32 unless specified otherwise.