### v2

# Controlled Low-Rank Adaptation with Subspace Regularization for Continued Training on Large Language Models

Abstract
Large language models (LLMs) exhibit remarkable capabilities in natural language processing but face catastrophic forgetting when learning new tasks, where adaptation to a new domain leads to a substantial decline in performance on previous tasks. In this paper, we propose Controlled LoRA (CLoRA), a subspace regularization method on LoRA structure. Aiming to reduce the scale of output change while introduce minimal constraint on model capacity, CLoRA imposes constraint on the direction of updating matrix’s null space. Experimental results on one-stage LLM finetuning tasks and continual learning settings highlight the superority of CLoRA as a effective parameter-efficient finetuning method with catastrophic forgetting mitigating. Further investigation for model parameters indicates that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
Controlled Low-Rank Adaptation with Subspace Regularization for Continued Training on Large Language Models
Yuheng Lu1, Bingshuo Qian1, Caixia Yuan1, Huixing Jiang2, Xiaojie Wang1,thanks: Corresponding author 1Beijing University of Posts and Telecommunications, 2LI Auto Inc. {yuheng.lu, bsqian, yuancx, xjwang}@bupt.edu.cn, jianghuixing@lixiang.com
1 Introduction
Large language models (LLMs) demonstrate remarkable capabilities in natural language tasks. However, when performing continued training on additional datasets, a key challenge may faced, known as catastrophic forgetting McCloskey and Cohen (1989), where adaptation to a new domain leads to a substantial decline in performance on previous tasks.
Existing approaches to mitigate catastrophic forgetting can be broadly categorized into data-based, architecture-based, and learning-based methods Wang et al. (2023a). Data-based methods de Masson D’Autume et al. (2019) primarily based on rehearsing prior training data, which raises data privacy concerns. Additionally, for LLMs, obtaining the necessary prior training data samples is challenging due to their training on massive datas. Architecture-based methods Wang et al. (2023c); Razdaibiedina et al. (2023) introduce isolated parameters for each continued training stage for reducing interference. In contrast, learning-based methods train in the shared vector space, controlling learning process by adding regularization terms to the loss or employing specific optimization designs. Inference for architecure-based methods typically involves a selection process Gurbuz and Dovrolis (2022); Kang et al. (2022), which is more complex than that for learning-based methods. As continued trained LLMs are generally regarded as foundation models, flexibility is essential for their broader applications. Consequently, due to deployment considerations, learning-based methods are preferred over architecture-based methods for LLMs.
The core idea of learning-based methods is to constrain parameter updates, which aligns precisely with the Parameter-Efficient FineTuning (PEFT) research paradigm of LLMs. Although initially proposed for computational efficiency, PEFTs have demonstrated to learn less and forget lessBiderman et al. (2024), primarily due to their constrained model capacity. Notably, a well-established insight that related to learning-based methods in PEFT research is that LLMs are primarily finetuned within a specific low-rank subspace, this insight has led to the development of the Low-Rank Adaptation method (LoRA)Hu et al. (2021).
However, LoRA imposes no restrictions on parameter updates beyond the low-rank constraint, and matrix perturbation theory suggests that even low-rank updates can significantly influence matrix properties Sherman (1949); Davis and Kahan (1970). For instance, in an extreme case, it is theoretically possible to learn a model that eliminates all top-k principal components (optimal rank-k approximation) through a rank-k update, thus destroy most of the base model’s ability. Therefore, LoRA would be benifit from more constraints for mitigating catastrophic forgetting. However, more constraints would reduce model capacity for updating, which influences the effectiveness of training. For instance, adding L2 regularization significantly restricts the norm of the updating matrix. Consequently, effective management of the capacity-forgetting balancing has become a major concern.
[…]
Figure 1: Illustration of the intuition behind our approach. For input $x$, the component in $\mathrm{Null}(\Delta W)$ (null space of the updating matrix $\Delta W$) would be ignored, the change of output $\Delta y$ is obtained only from the component in $\mathrm{Row}(\Delta W)$ (row space of $\Delta W$, the orthogonal complement of $\mathrm{Null}(\Delta W)$). CLoRA introduces a pre-defined subset of $\mathrm{Null}(\Delta W)$ by imposing orthogonal regularization with pre-defined matrix $P$.
To address this concern, in this work, we propose Controlled LoRA (CLoRA), a subspace regularization method on LoRA structure. We start the design of CLoRA from the perspective of the null space of updating matrix. The intuition behind CLoRA is illustrated in Figure 1, where the output change $\Delta y$ is derived from applying the updating matrix $\Delta W$ on the component of the input $x$ that falls within the row space of $\Delta W$, while components in the null space are ignored. Under this intuition, for reducing the scale of output change, options include reducing the scale of $\Delta W$, and encouraging more input component fall in the null space of $\Delta W$. The former is more related to model capacity, and for concerns of capacity-forgetting balancing, we focus on the latter.
The dimension of the null space for the updating matrix is directly determined by the rank of it, which LoRA already addressed. A key factor remains, the direction of null space, which influence input components that fall in, but free-learned LoRA does not constraint. CLoRA constraint the direction of null space of updating matrix by introducing a pre-defined subspace, this is implemented by orthogonal regularization with a pre-defined matrix. Unlike methods that impose restrictions on rank or norm, which significantly influence model capacity, CLoRA introduces constraint on the direction of the null space. We take experiments on commonly used one-stage LLM finetuning evaluations and continual learning evaluations, results indicate the superiority of CLoRA as an effective approach for parameter-efficient finetuning with catastrophic forgetting mitigating. Additionally, we take analysis on parameters of the learned model, results show that CLoRA reduces the scale of output change with minimal impact on model capacity.
Our contributions are summarized as follows,
  • We propose CLoRA, a subspace regularization method on LoRA, which serves as an advanced parameter-efficient finetuning technique with catastrophic forgetting mitigating for LLMs.
  • Our proposed CLoRA demonstrates superior performance on both in-domain and out-domain evaluation in commonly used one-stage LLM finetuning setting. Additionally, it showns remarkable mitigating of catastrophic forgetting in continual learning setting.
  • Parameter investigation results indicate that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
2 Related Works
2.1 Mitigating Catastrophic Forgeting
Catastrophic forgetting is a significant challenge in various transfer learning scenarios, including continual learning Wang et al. (2023a) and LLM finetuning Wu et al. (2024). In these settings, continued training on new tasks may impair abilities of the pre-trained model. Approaches for mitigating catastrophic forgetting can be broadly categorized into data-based, architecture-based and learning-based methods.
Data-based methods
primarily based on rehearsal of prior training data or representation, de Masson D’Autume et al. (2019) introduce an episodic memory for experience rehearsal, Rebuffi et al. (2017); Chaudhry et al. (2019) selects previous training data for rehearsaling. For LLMs, acquiring the necessary prior training data is challenging due to the extensive amount of data used in their training. Instead, the concept of rehearsal is commonly adopted by mixing data from general domains for LLM continued training. This approach is generally orthogonal to model-related methods, thus we will not discuss it further.
Architecture-based methods
Wang et al. (2023c); Razdaibiedina et al. (2023) introduce isolated parameters for each continued training stage to reduce interference. Wang et al. (2023c) use isolated parameters for each task, and enables a selecting mechenism during inference. Progressive Prompts Razdaibiedina et al. (2023) sequentially concatenates prompts for each task with previously learned prompts. These architecture-based methods generally require specific techniques for inference and continued training, resulting in a lack of flexibility, particularly in the context of LLMs.
Learning-based methods
performs continued training in a shared vector space, controlling learning process by adding regularization term on loss or applying specific optimization designs. Notabley, O-LoRA Wang et al. (2023b) introduce regularization with previous continual learned parameters for reducing interference in the multi-stage training setting. Our proposed CLoRA imposes orthogonal regularization similar to O-LoRA, but the regularization matrix is not restricted to be the previous learned parameter, thus CLoRA can be used for one-stage continued training whereas O-LoRA not.
2.2 LoRA and Subspace Tuning
Parameter-Efficient FineTuning (PEFT) Han et al. (2024) aims to tune models with minimizing computational resources, which is widely used for large-scale models including LLMs. Among these methods, LoRA Hu et al. (2021) and its subsequent variants Wang et al. (2024a); Liu et al. (2024) learn a low-rank decomposition for updating parameter matrices, and could be categorized into learning-based continued training method, which is the focus of our work.
The core insight of LoRA is to tune model within a low-rank subspace, and with no additional constraints imposed on this tuning subspace. Some subsequent works delve deeper into the tuning subspace to mitigate catastrophic forgetting for LLM continued training, MiLoRA Wang et al. (2024a) and PiSSA Meng et al. (2024) use singular value decomposition (SVD) components of the original parameters for LoRA initialization, with MiLoRA uses minor components while PiSSA uses major components; O-LoRA Wang et al. (2023b) introduce orthogonal regularization for each LoRA subspace. Our proposed CLoRA also falls within this category, differing from the selection and utilization of the focused subspace.
[…]
Figure 2: Illustration of CLoRA on typical decoder-only transformer based LLMs. LoRA updating is applied on v-proj in multi-head attention layer for each layer. CLoRA add orthogonal loss computes from trainable LoRA parameters (A and B) to the original language modeling loss.
3 Prelimilaries
Notation Description
$W$ parameter matrix in base model
$\Delta W$ updating of the parameter
$x$ input for $W$
$y$ output for $W$, $y=Wx$
$\Delta y$ output change, $\Delta y=\Delta Wx$
$||v||$ L2 norm of vector $v$
$||A||$ L2 norm(largest singular value)
of matrix $A$
$r$ rank of updating matrix
$k$ number of regularization vectors
Table 1: Notations.
3.1 Notations
The notations commonly used in this paper are summarized in table 1. We provide some additional notes here. While generally used for denote the input and output of the while model, we denote $x,y$ as input and output to a single linear layer, represented by $W$. $||A||$ denotes L2 norm (largest singular value) in our paper, instead of Frobenius norm ($||A||_{F}=\sqrt{\sum A[i,j]^{2}}$). $r$ and $k$ are most important hyperparameters for CLoRA, $r$ is the rank of updating matrix, which is used in all LoRA works, $k$ is the number of regularization vectors(column of regularization matrix) in CLoRA.
3.2 Problem Definition
Catastrophic forgetting menifest as performance decline on tasks from previous domain when training on new domain. In this work, we aim to mitigate catastrophic forgetting in LLM finetuning and continual learning settings.
3.2.1 LLM Finetuning
In this setting, we conduct experiments on one-stage LLM finetuning, To evaluate this, we conduct both in-domain tasks (demonstrating the effectiveness of training) and out-domain tasks (from previous domain, indicating the degree of forgetting) for LLM finetuning. Specifically, we finetune a base LLM on one training dataset, then take in-domain and out-domain evaluations. Note that there is no clear domain specific for base LLMs, but benchmarks exist for evaluating the ability of LLMs on wide range of domains Gao et al. (2024), and we take those with minimal overlap with training data for out-domain evaluation.
3.2.2 Continual Learning
Continual learning focuses on developing learning algorithms to accumulate knowledge on non-stationary dataWang et al. (2023b). In this setting, we conduct experiments for multi-stage finetuning. Specifically, we finetune the model on a sequence of tasks $D_{1},\dots,D_{t}$, where each task $D_{t}$ contains a pair of train and test datasets $D_{t}=(D^{train}_{t},D^{test}_{t})$. The $t$-th model with finetuned sequentially on $D^{train}_{1},\dots,D^{train}_{t-1}$ is tested over all previous test datasets $D^{test}_{1},\dots,D^{test}_{t-1}$.
4 Method
In this section, we introduce Controlled Low-Rank Adaptation (CLoRA) method. We illustrate the application of CLoRA in transformer-based LLMs in Figure 2. CLoRA shares the same modeling structure with LoRA, but imposes on orthogonal regularization term computed using LoRA parameters into the loss function.
CLoRA Modeling
Consistent with LoRA, CLoRA decomposes the updating for a parameter matrix $W$ to a multiplication of two lor-rank matrices $\Delta W=AB^{T}$, where $W,\Delta W\in\mathbb{R}^{m\times n}$, $A\in\mathbb{R}^{m\times r}$, $B\in\mathbb{R}^{n\times r}$, $r\ll m,n$.
CLoRA computes orthogonal regularization for $A$ and $B^{T}$ with untrainable pre-defined matrix $P_{A}\in\mathbb{R}^{m\times k}$ and $P_{B}\in\mathbb{R}^{n\times k}$, where $k$ is a hyperparameter controlling the size of regularization matrix, larger $k$ introduces more constraint. The orthogonal regularization loss on one LoRA parameter $A$ is defined as
$L_{orth}(A,P_{A})=\sum_{i,j}||AP_{A}^{T}[i,j]||^{2}$ (1)
where $A\in\mathbb{R}^{m\times r}$, $P_{A}\in\mathbb{R}^{m\times k}$. $L_{orth}(A,P_{A})$ regularize on orthogonality of every $(A[:,i],P_{A}[:,j])$ pairs. The final loss of CLoRA in a transformer-based LLM is defined as
$L_{LM}(\Theta,input)+$ (2)
$\lambda\sum_{i}(L_{orth}(A_{i},P_{A_{i}})+L_{orth}(B_{i}^{T},P_{B_{i}}^{T}))$
where $L_{LM}(\Theta,input)$ is the original language model loss on text input and LLM parameters $\Theta$, the summation on $L_{orth}$ is over index of all trainable parameter matrices. $\lambda$ controls the weighting of orthogonal loss, we set it to 1 as default.
Initialization
Following LoRAHu et al. (2021), we initialize $A$ with gaussian noise and $B$ is zeros, ensuring $\Delta W$ is zero at the begining of training.
For the CLoRA regularization matrices, following the priciple of Occam’s Razor, we adopt the most simple random initialization here. For uniform regularization over each row in regularization matrices, we suggest using orthogonal initialization. Specifically, for regularization matrix $P\in\mathbb{R}^{m\times k}$, $||P[:,i]||=1$ for every $i$, and $P[:,i]P[:,j]=0$ for $i\neq j$.
5 Experiments and Analysis
5.1 One-Stage LLM Finetuning
In this section, we conduct experiments on one-stage LLM finetuning to evaluate our proposed CLoRA as a parameter-efficient finetuning method. We aim to answer the following research questions,
  • RQ1: Does CLoRA performs effectively as a parameter-efficient finetuning method for LLMs with catastrophic forgetting mitigating?
  • RQ2: How the size of regularization matrix $k$ influence the performance of CLoRA? Does it differs across tasks?
  • RQ3: How does CLoRA demonstrate superiority on capability-forgetting balancing?
5.1.1 Datasets and Tasks
Following previous works on PEFTLiu et al. (2024); Wang et al. (2024a), we conduct experiments on commonsense reasoning tasks and math tasks.
Commonsense Reasoning Setting
We use Commonsense170K Hu et al. (2023) for finetuning. For in-domain evaluation, eight commonsense reasoning datasets are used, including BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2019), ARC-e, ARC-c Clark et al. (2018), and OBQA Mihaylov et al. (2018). The tasks are formulated as multiple-choice problem, and we report accuracy based on the last checkpoint.
For out-domain evaluations, BIG-Bench-Hard Suzgun et al. (2022) and MMLU-Pro Wang et al. (2024b) are used. These benchmarks encompass challenging subsets of tasks across a wide range of domains and are widely employed for evaluating the capabilities of LLMs. Additionally, they include samples that are more complex than those in our training data, ensuring minimal overlap. We use lm-eval Gao et al. (2024), available with MIT License, for reporting out-domain evaluation.
Math Setting
We use MetaMathQA Yu et al. (2024) for finetuning, which contains 395K samples augmented from the training set of GSM8K Cobbe et al. (2021) and MATHHendrycks et al. (2021). We use test set of GSM8K and MATH for evaluation and report the results on the last checkpoint.
5.1.2 Comparison Methods
  • LoRA Hu et al. (2021) is a widely-used parameter-efficient finetuning technique, and it serves as the foundation of our proposed CLoRA.
  • DoRA Liu et al. (2024) is a recent work on structure improvement of LoRA, we include it as a baseline for improved LoRA.
  • PiSSA Meng et al. (2024) and MiLoRA Wang et al. (2024a) are two variants of LoRA, both employing SVD components for LoRA initialization, MiLoRA use minor components while PiSSA use major. Notably, MiLoRA can be categorized as a catastrophic forgetting mitigating method.
  • Reducing the updating rank(-r*): Lower rank $r$ imposes stricter constraints on the updating matrix. We maintain a consistent rank across all methods and consider variations in rank as a separate baseline.
  • L2 regularization(-L2) introduces L2 regularization for trainable parameters, serving as a fundamental approach to limit updates.
  • CLoRA: Our proposed CLoRA method, with random initialized regularization matrix.
In-domain Out-domain
Method BQ PQ SQ HS WG ACe ACc OQ Avg. BBH MMLU
LLaMA2-7b - - - - - - - - - 34.91 18.56
LoRA 71.9 80.9 78.9 90.3 83.5 83.0 70.2 80.8 79.9 26.69 14.46
DoRA 73.0 81.9 80.3 90.2 82.8 84.6 69.4 81.8 80.5 28.24 11.67
PiSSA 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8 29.54 11.33
MiLoRA 71.5 82.0 80.0 91.0 83.0 82.3 68.9 81.2 80.0 25.14 17.74
LoRA-r8 71.0 80.5 78.1 90.0 83.0 81.1 68.5 78.0 78.8 26.90 14.58
LoRA-r16 71.0 81.8 78.9 90.3 81.1 83.1 69.7 82.2 79.8 26.73 11.54
LoRA-L2 70.3 83.0 80.2 92.7 83.1 84.2 71.2 81.4 80.8 32.93 16.59
CLoRA-k128 72.7 84.1 77.7 91.6 83.0 85.3 69.9 81.6 80.7 30.82 12.07
CLoRA-k256 71.3 83.2 79.1 92.4 83.2 84.5 71.0 81.0 80.7 31.92 17.81
CLoRA-k512 72.8 83.0 79.5 93.0 83.9 85.7 73.0 84.8 82.0 34.32 17.00
CLoRA-k1024 73.3 84.8 79.6 91.1 86.1 86.9 73.1 85.6 82.6 36.49 19.52
CLoRA-k2048 73.7 84.5 80.9 94.5 85.9 88.1 75.9 86.0 83.7 38.67 20.59
Table 2: Results for our proposed CLoRA and baselines for in-domain commonsense reasoning evaluations and out-domain LLM benchmarks, with accuracy scores (%) reported. Bold font indicates the highest performance for each task across all compared PEFT methods.
Method GSM8K MATH
LoRA 60.58 16.88
PiSSA 58.23 15.84
MiLoRA 63.53 17.76
CLoRA-k64 64.29 17.52
CLoRA-k128 64.59 18.38
CLoRA-k256 63.45 17.58
Table 3: Math evaluation on GSM8K and MATH, with accuracy scores (%) reported.
5.1.3 Experimental Configuration
We use the same base LLM choice LLaMA-2-7B Touvron et al. (2023) and hyperparameter configurations as Wang et al. (2024a). Details are listed in Appendix A. Notably, we use 32 (commonsense reasoning) and 64 (math) for updating matrix rank $r$ as default for all methods if not explicitly specified. For the size of CLoRA regularization matrix, we select $k$ in $[128,256,512,1024,2048]$ for commonsense reasoning and $[64,128,256]$ for more challenging math setting. For LoRA-L2, 1e-5 is used for weighting of L2 regularization, we note that 1e-4 is also tested, but too large for getting effective finetuning. We report results finetuned on LLaMA-2-7B here, more results are listed in Appendix A.
5.1.4 Main Results (RQ1)
For commonsense reasoning setting, we report the results of in-domain evaluation and out-domain LLM benchmarks in Table 2. The results demonstrate that CLoRA outperforms on all datasets, surpassing the best baseline for in-domain evaluation by an average accuracy of 2.9 points. Results for math setting also demonstrate the superority of CLoRA over previous LoRA baselines (Table 3).
These outcomes suggest that, although primarily proposed for mitigating catastrophic forgetting, CLoRA also serves as an effective PEFT method. We attribute this to the nature of LLM finetuning, which is an instance of transfer learning. The performance of LLM finetuning is strongly correlated with the base model’s ability, when catastrophic forgetting occurs during training, the base model’s strength may diminish. Therefore, we claim that a method with effective capacity-forgetting balancing would exhibit strong effectiveness in LLM finetuning.
For out-domain evaluation, results show that all baselines underperform the base model, highlighting the severe issue of catastrophic forgetting in this setup. Notably, our proposed CLoRA not only outperforms all baselines by a significant margin but also surpasses the base model’s performance. We attribute this to CLoRA’s effective capacity-forgetting balancing, which enables the extraction of generally useful knowledge from the commonsense reasoning training dataset.
The superior performance in both in-domain and out-domain evaluations demonstrates that CLoRA serves as an effective parameter-efficient finetuning method with catastrophic forgetting mitigation.
The superior performance in both in-domain and out-domain evaluations demonstrates that our proposed CLoRA serves effectively as a parameter-efficient finetuning method with catastrophic forgetting mitigating. Thus, we answer RQ1.
5.1.5 Evaluating for different CLoRA $k$ (RQ2)
The size of the regularization matrix $k$ is a crucial hyperparameter in CLoRA, balancing the trade-off between model capacity and the degree of forgetting. We focus here on how $k$ influence the performance of finetuning LLM with CLoRA, and investigate whether the optimal $k$ is consistent across tasks.
In commonsense reasoning setting, results show that larger $k$ leads to better performance in both in-domain and out-domain evaluations (Table 2). In math setting, unlike the upward trend in commonsense reasoning setting, performance decreases when $k$ exceeds 128(Table 3). We attribute this discrepancy to the complexity of math tasks, which require greater model capacity during finetuning.
Emperical results support the intuitive claim that larger $k$ imposes more restrictions on updates, which helps mitigating catastrophic forgetting but potentially limiting finetuning model capacity and harming performance.
Thus, we answer RQ2 by demonstrating that the optimal $k$ depends on task complexity. Notably, our proposed CLoRA provides flexibility in balancing capacity and forgetting by adjusting $k$, we suggest choosing a smaller $k$ for more challenging tasks.
5.1.6 Understanding Capacity-Forgetting Balancing(RQ3)
To answer RQ3, we investigate the parameter of trained models to quantify the capacity-forgetting balancing issue.
Method $||\Delta W||$ $\mathbb{F}$
reference 2.42
LoRA 22.63 0.79
MiLoRA 24.32 0.92
LoRA-r16 12.70 1.03
LoRA-r8 6.45 0.95
LoRA-L2 2.07 0.29
CLoRA-k128 10.84 0.36
CLoRA-k256 10.25 0.34
CLoRA-k512 8.19 0.27
CLoRA-k1024 6.64 0.21
CLoRA-k2048 5.00 0.14
Table 4: Measuring model updating capacity($||\Delta W||$, larger for more capacity) and degree of forgetting($\mathbb{F}$, lower for less forgetting) for trained models.
Measuring Model Capacity and Degree of Forgetting
Consider that catastrophic forgetting primarily arises from output changes caused by parameter updating, the greater the impact of these updates, the more severe the catastrophic forgetting may be. We measure the degree of forgetting with the relative scale of output change in the parameter level, to be specific, for updating matrix $\Delta W$, with input $x$, the relative scale of output change (denoted as $\mathbb{F}$) is defined as
$\mathbb{F}(\Delta W,x)=\frac{||\Delta Wx||}{||x||}$ (3)
We highlight the role of $x$ in the measurement of $\mathbb{F}$, as it reflects the real world case. Specifically, we sample 100 text data from test set and collect input $x$ for each parameter from the model forward pass.
To measure model capacity, we note that there is a gap between theoretical capacity of a model Abu-Mostafa (1989) and the practical outcome of the learned model. Therefore, we delegate the measurement of model capacity to the scale of the parameters in the learned model. Specifically, we measure the L2 norm $||\Delta W||$ for each updating parameter matrix.
Results and Analysis
We report the measurements averaged over all tokens and all updating parameters in Table 4. All models use LoRA rank $r$ of 32 unless specified otherwise.
The “reference” row is computed using the LoRA trained model, noting the output scale of original parameter $W$ instead of $\Delta W$. Compared with “reference” and LoRA, the difference of $\mathbb{F}$ is not far, suggesting that LoRA training indeed introduces significant output change, thus still prone to catastrophic forgetting.
For MiLoRA, although intuitively promising, without effective control during training, it did not mitigate catastrophic forgetting, as evidenced by both downstream evaluations (Table 2) and the similar $\mathbb{F}$ and $\Delta W$ with LoRA.
For LoRA with lower rank (r8/16), after training, with $||\Delta W||$ indicates the reduction of capacity, $\mathbb{F}$ does not show a decrease. Although theoretically, reducing the rank of the update matrix can increase the dimension of the null space and help to reduce the scale of output change, results not show this case. This suggests that altering $r$ may not a effctive way to alter forgetting.
For LoRA-L2, $\mathbb{F}$ indicates that it indeed mitigate forgetting, but in a large cost of capacity, demonstrated by the very small $||\Delta W||$.
For our proposed CLoRA, $\mathbb{F}$ shows a significantly reduce the scale of output change, while a relatively larger $||\Delta W||$ is maintained. This indicates that CLoRA minimizes catastrophic forgetting caused by large updates while having a subtle impact on model capacity. Thus we answer RQ3 that CLoRA performs effectively on capacity-forgetting balancing.
Standard CL Benchmark Large Number of Tasks
Method Order-1 Order-2 Order-3 avg. Order-4 Order-5 Order-6 avg.
SeqFT 18.9 24.9 41.7 28.5 7.4 7.4 7.5 7.4
SeqLoRA 44.6 32.7 53.7 43.7 2.3 0.6 1.9 1.6
IncLoRA 66 64.9 68.3 66.4 63.3 58.5 61.7 61.2
Replay 55.2 56.9 61.3 57.8 55 54.6 53.1 54.2
EWC 48.7 47.7 54.5 50.3 45.3 44.5 45.6 45.1
LwF 54.4 53.1 49.6 52.3 50.1 43.1 47.4 46.9
L2P 60.3 61.7 61.1 60.7 57.5 53.8 56.9 56.1
LFPT5 67.6 72.6 77.9 72.7 70.4 68.2 69.1 69.2
O-LoRA 75.4 75.7 76.3 75.8 72.3 64.8 71.6 69.6
CLoRA 79.7 79.1 78.2 79.0 70.7 65.6 68.2 68.1
PerTaskFT 70.0 70.0 70.0 70.0 78.1 78.1 78.1 78.1
MTL 80.0 80.0 80.0 80.0 76.5 76.5 76.5 76.5
Table 5: Results on two CL benchmarks with T5-large base model. Averaged accuracy after training on the last task is reported. Bold font indicates the highest performance across all compared CL methods.
5.2 Continual Learning
5.2.1 Experimental Setup
To demonstrate the effectiveness of CLoRA for continual learning(CL) setting, we conduct experiments on standard CL benchmark and more challenging large number of tasks benchmark, following the experiment setup of O-LoRAWang et al. (2023b).
Datasets and Tasks
The standard CL benchmark consists of five text classification datasetsZhang et al. (2015). The large number of tasks benchmark consists of 15 datasets Razdaibiedina et al. (2023), include tasks for natural language understanding and text classification. Task samples follows previous work Wang et al. (2023b). Details for tasks are listed in Appendix B.
Comparison Methods
We compare CLoRA with normal finetuning baselines and previous CL methods. We include non CL results that train separate model for each task (PerTaskFT) and multi-task learning (MTL) as reference.
  • Normal Finetuing baselines include sequentially training on same parameter space with full parameter finetune (SeqFT) and LoRA (SeqLoRA), and incremental learning of new LoRA parameters on a sequential series of tasks.
  • Continual Learning methods include data-based methods Replay; architecture-based methods L2PWang et al. (2022), LFPT5Qin and Joty (2022)O-LoRAWang et al. (2023b); and learning-based methods EWCKirkpatrick et al. (2017), LwFLeibe et al. (2016). Details for these methods are listed in Appendix B.
Experimental Configuration
Following O-LoRAWang et al. (2023b), we use T5-large as base model, and finetune on each task with specified order(Appendix B). We train each task with one epoch, with constant learning rate of 1e-3, batch size of 64, dropout rate of 0.1, weight decay rate of 0, and LoRA dim $r$ of 8. CLoRA regularization matrix size $k$ is set to 256.
5.2.2 Results and Analysis
We report the results in Table 5. Results demonstrate that CLoRA outperforms all comparision methods, include the most related strong baseline O-LoRA, with a notable margin in the stadard CL benchmark. We attribute this to the advantage of CLoRA toward O-LoRA: 1. CLoRA helps learning in the first finetuning stage while O-LoRA not; 2. CLoRA can independently alter $k$ for balancing learning and forgetting, while “$k$” equivalent in O-LoRA is restrained by LoRA $r$.
In the large number of tasks benchmark, CLoRA performs near the strong baseline O-LoRA and LFPT5. We note that vanilla CLoRA with random regularization matrix is a learning-based method, without machanism for isolating finetuning tasks.
6 Conclusion
In this paper, we introduce Controlled Low-Rank Adaptation(CLoRA), a simple yet effective parameter-efficient finetuning method for LLMs that mitigates catastrophic forgetting. We investigate the effectiveness of CLoRA on both one-stage LLM finetuning and continual learning settings. Experiment results demonstrate the effectiveness of CLoRA as a parameter-efficient finetuning method with catastrophic forgetting mitigating. Further investigation for model parameters indicates that CLoRA effectively balances the trade-off between model capacity and degree of forgetting.
7 Limitations
There are still several limitations that we reserve for future work: 1) We use the simplest random initialization for regularization matrix, insight for more dedicated choice may benifit CLoRA learning, such as combine CLoRA with architecture-based continual learning method(O-LoRA). 2) We delegate the measurement of model capacity and degree of forgetting to simple measurement of scale. Although these measurements reveal significant differences between CLoRA and previous works, we believe that further investigation would aid in the design of methods with stronger capacity-forgetting balancing capability.
Appendix A Experiment Details for One-Stage Finetuning
A.1 Hyperparameter Settings
Table 6 shows our detailed hyperparameters. This setting follows MiLoRAWang et al. (2024a) and DoRALiu et al. (2024).
A.2 Computation Environment
All of our experiments are conducted on 8 NVIDIA A800 GPUs. All methods for LoRA subsequents use Huggingface peft library [^1] https://github.com/huggingface/peft, training is conducted using trainer in Huggingface transformers library [^2] https://github.com/huggingface/transformers, with DeepSpeed ZeRORajbhandari et al. (2020) intergration.
A.3 Additional CLoRA variants
We use the simplest random initialization for CLoRA regularization matrix in the main paper. Considering the idea of PiSSA and MiLoRA that explore the roles of singular value decomposition (SVD) components in LLM parameters, we adopt this intuition to initialize CLoRA regularization matrices from SVD. For a SVD decomposition of parameter $W=USV^{T}$ with rank $r$, where $W\in\mathbb{R}^{m\times n}$, $U\in\mathbb{R}^{m\times r}$, $S\in\mathbb{R}^{r\times r}$ is a diagonal matrix, $V\in\mathbb{R}^{n\times r}$. For CLoRA updating $\Delta W=AB^{T}$, we initialize the regularization matrix $P_{A}\in\mathbb{R}^{m\times k}$ as $U[:,\mathbf{s}]$, $P_{B}\in\mathbb{R}^{n\times k}$ as $V[:,\mathbf{s}]$, where $\mathbf{s}$ is a list of selecting index with length $k$. We add two CLoRA variants as follows, and conduct experiments on commonsense reasoning setting,
  • CLoRA-major: Use SVD major components to initialize CLoRA regularization matrix.
  • CLoRA-minor: Use SVD minor components to initialize CLoRA regularization matrix.
A.4 Full Results on Commonsense Finetuning
Model PEFT BoolQ PIQA SIQA HS WG ARC-e ARC-c OBQA Avg.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA-2-7B LoRA 71.9 80.9 78.9 90.3 83.5 83.0 70.2 80.8 79.9
PiSSA 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA 71.5 82.0 80.0 91.0 83.0 82.3 68.9 81.2 80.0
DoRA 73.0 81.9 80.3 90.2 82.8 84.6 69.4 81.8 80.5
LoRA-r8 71.0 80.5 78.1 90.0 83.0 81.1 68.5 78.0 78.8
LoRA-r16 71.0 81.8 78.9 90.3 81.1 83.1 69.7 82.2 79.8
LoRA-L2-0.0001 - - - - - - - - -
LoRA-L2-0.00001 70.3 83.0 80.2 92.7 83.1 84.2 71.2 81.4 80.8
CLoRA-random-k128 72.7 84.1 77.7 91.6 83.0 85.3 69.9 81.6 80.7
CLoRA-random-k256 71.3 83.2 79.1 92.4 83.2 84.5 71.0 81.0 80.7
CLoRA-random-k512 72.8 83.0 79.5 93.0 83.9 85.7 73.0 84.8 82.0
CLoRA-random-k1024 73.3 84.8 79.6 91.1 86.1 86.9 73.1 85.6 82.6
CLoRA-random-k2048 73.7 84.5 80.9 94.5 85.9 88.1 75.9 86.0 83.7
CLoRA-major-k128 72.4 81.9 77.9 83.9 82.4 84.4 70.0 82.6 79.4
CLoRA-major-k256 73.2 83.5 79.6 93.0 83.3 88.1 72.6 84.2 82.2
CLoRA-major-k512 73.6 83.7 79.9 93.4 83.9 86.4 73.0 86.0 82.5
CLoRA-major-k1024 73.2 85.5 80.5 94.3 85.7 87.2 75.9 85.4 83.5
CLoRA-major-k2048 73.9 84.8 80.6 95.0 85.3 87.7 76.5 84.6 83.6
CLoRA-minor-k128 71.5 82.7 78.7 91.8 83.2 85.0 70.9 81.6 80.7
CLoRA-minor-k256 72.6 83.5 80.2 91.3 85.4 85.4 72.1 83.6 81.8
CLoRA-minor-k512 73.0 84.0 80.1 93.1 82.0 86.4 72.9 84.4 82.0
CLoRA-minor-k1024 73.1 83.7 79.2 93.7 84.8 87.1 73.2 83.2 82.3
CLoRA-minor-k2048 72.9 84.2 80.8 93.7 85.3 87.2 73.5 86.0 83.0
LLaMA-3-8B LoRA 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
PiSSA 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
DoRA 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8 85.2
CLoRA-random-k128 75.5 89.1 81.6 95.9 87.9 92.6 81.8 86.8 86.4
CLoRA-random-k256 75.3 88.8 81.4 85.7 88.7 92.7 82.3 88.4 85.4
CLoRA-random-k512 75.9 89.3 82.6 96.3 88.9 92.1 82.9 86.8 86.9
CLoRA-random-k1024 76.5 89.1 82.1 96.3 88.6 93.0 81.7 90.0 87.2
CLoRA-random-k2048 76.2 90.0 82.7 96.6 88.8 93.3 83.4 89.2 87.5
Table 7: In-domain results on commonsense reasoning evaluations, with accuracy scores (%) reported. Bold font indicates the highest performance for each dataset across the different PEFT methods for each base model.
Model PEFT BBH MMLU-Pro Avg.
LLaMA-2-7B - 34.91 18.56 26.74
LoRA 26.69 14.46 20.58
PiSSA 29.54 11.33 20.44
MiLoRA 25.14 17.74 21.44
DoRA 28.24 11.67 19.96
LoRA-r8 26.90 14.58 20.74
LoRA-r16 26.73 11.54 19.13
LoRA-L2-0.00001 32.93 16.59 24.76
CLoRA-random-k128 30.82 12.07 21.45
CLoRA-random-k256 31.92 17.81 24.87
CLoRA-random-k512 34.32 17.00 25.66
CLoRA-random-k1024 36.49 19.52 28.01
CLoRA-random-k2048 38.67 20.59 29.63
CLoRA-major-k128 32.69 18.09 25.39
CLoRA-major-k256 35.11 18.89 27.00
CLoRA-major-k512 35.81 19.88 27.85
CLoRA-major-k1024 37.06 19.73 28.40
CLoRA-major-k2048 38.83 20.08 29.46
CLoRA-minor-k128 34.06 17.03 25.55
CLoRA-minor-k256 33.16 17.11 25.13
CLoRA-minor-k512 35.42 18.97 27.20
CLoRA-minor-k1024 37.08 18.87 27.98
CLoRA-minor-k2048 40.96 20.37 30.67
Table 8: Out-domain results on two LLM benchmarks, with accuracy scores (%) reported. Bold font indicates the highest performance for each benchmark across all methods.
We report the full results that we conducted on in-domain evaluation(Table 7) and out-domain evaluation(Table 8) for commonsense reasoning finetuning. Results for LLaMA-3-8b are also included for CLoRA-random. All models use LoRA rank $r$ of 32 unless specified otherwise.
A.4.1 Analysis for different CLoRA variants
Results indicate that the choice of regularization matrix does influence the effectiveness of CLoRA, albeit not significantly. Generally, we recommend using random initialization (CLoRA-random) or initialization from major SVD components (CLoRA-major).
Appendix B Detailed Experiment Setups for Continual Learning
Dataset name Category Task Domain
Yelp CL Benchmark sentiment analysis Yelp reviews
Amazon CL Benchmark sentiment analysis Amazon reviews
DBpedia CL Benchmark topic classification Wikipedia
Yahoo CL Benchmark topic classification Yahoo Q&A
AG News CL Benchmark topic classification news
MNLI GLUE NLI various
QQP GLUE paragraph detection Quora
RTE GLUE NLI news, Wikipedia
SST-2 GLUE sentiment analysis movie reviews
WiC SuperGLUE word sense disambiguation lexical databases
CB SuperGLUE NLI various
COPA SuperGLUE QA blogs, encyclopedia
BoolQA SuperGLUE boolean QA Wikipedia
MultiRC SuperGLUE QA various
IMDB SuperGLUE sentiment analysis movie reviews
Table 9: Summary of datasets used in the continual learning setting.
Order Task Sequence
1 dbpedia $\rightarrow$ amazon $\rightarrow$ yahoo $\rightarrow$ ag
2 dbpedia $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ yahoo
3 yahoo $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ dbpedia
4 mnli $\rightarrow$ cb $\rightarrow$ wic $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ boolqa $\rightarrow$ rte $\rightarrow$ imdb
$\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ multirc $\rightarrow$ yahoo
5 multirc $\rightarrow$ boolqa $\rightarrow$ wic $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte
$\rightarrow$ imdb $\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ yahoo
6 yelp $\rightarrow$ amazon $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb
$\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ multirc $\rightarrow$ boolqa $\rightarrow$ wic
Table 10: Order of finetuning in the continual learning setting.
B.1 Dataset Details
We list the details of the datasets used in Table 9. Order of finetuning are listed in Table 10.
B.2 Computation Environment
All of our experiments are conducted on 1 NVIDIA GeForce RTX 3090 GPU. All methods for LoRA subsequents use Huggingface peft library, training is conducted using trainer in Huggingface transformers library, with DeepSpeed ZeRO intergration.
B.3 Comparision Methods
Here we provide details for continual learning baselines for our continual learning experiment setting.
  • Replay: data-based method that replay samples from old tasks when learning new tasks to avoid forgetting.
  • L2P: architecture-based method that uses the input to dynamically select and update prompts from the prompt pool in an instance-wise fashion.
  • LFPT5: architecture-based method that continuously train a soft prompt that simultaneously learns to solve the tasks and generate training samples for replay.
  • EWC: learning-based method that finetune the whole model with a regularization loss that prevents updating parameters that could interfere with previously learned tasks.
  • LwF: learning-based method that constrains the shared representation layer to be similar to its original state before learning the new task.
  • O-LoRA: architecture and learning-based method that prevent subsequent LoRA update interfere previous.